Two forces P and 2P acting at a point have a resultant R which is perpendicular to the line P, find the angle between the forces.

Grade:12th pass

1 Answers

Eshan
5 years ago
The resultant of vectors\vec{A}and\vec{B}(with angle\thetabetween them) forms an angle\alphawith\vec{A}such that

tan\alpha=\dfrac{Bsin\theta}{A+Bcos\theta}

Since\alpha=90^{\circ},

A+Bcos\theta=0\implies cos\theta=-\dfrac{A}{B}=-\dfrac{P}{2P}=-\dfrac{1}{2}

\implies \theta=120^{\circ}